# HeurekaBench: A Benchmarking Framework for AI Co-scientist

### v1

@@ 1 Introduction
Post-training algorithms for large language models (LLMs) with reinforcement learning (RL) have led to rapid progress in improving their reasoning capabilities (rafailov2023direct; shao2024deepseekmath; chai2025empowering). As a result, LLM-based agents have been developed to solve complex tasks through multi-turn interactions with an external environment (huang2022language; yao2023react; yang2024sweagent). These systems can be broadly characterized by a loop consisting of three stages: in the think stage, the agent needs to understand the user request and create a plan; in the act stage it generates an action depending on the plan; and in the observe stage it updates the context with the feedback of action execution for the next cycle.
Agentic systems offer promise for advancing scientific discovery by actively generating new hypotheses, designing and evaluating experiments, and iteratively improving through feedback (m2024augmenting; ghareeb2025robin; saeedi2025astroagents; huang2025automated). In this capacity, they could act as AI co-scientists, complementing human expertise and accelerating the scientific discovery process (yamada2025ai; gottweis2025towards). One scientific domain that has recently seen an influx of agents is single-cell biology (xiao2024cellagent; roohani2025biodiscoveryagent; alber2025cellvoyager; mitchener2025bixbench; huang2025biomni). A typical real-world use case in this domain involves autonomously proposing novel hypotheses, performing complex analyses of single-cell RNA-seq experimental datasets and delivering scientifically meaningful findings.
@@
Most existing data-driven scientific discovery benchmarks test task instruction following and answering a single computational question from experimental data (e.g., How many miRNAs remain significant at $p\leq 0.05$ after Benjamini-Hochberg correction? from mitchener2025bixbench and Train a VAE model and perform a 1-vs-all differential expression test for each cell type from chen2025scienceagentbench). However, a co-scientist should autonomously plan these questions as sub-steps within the workflow, rather than explicit user instructions. The closest to our work is BaisBench (luo2025benchmarking); however, it relies solely on the capabilities of LLMs to formulate questions from scientific studies, leading to invalidated questions that may not be possible to answer accurately. Thus, existing benchmarks lack an appropriate framework for evaluating open-ended, free-form, data-driven scientific responses that are required for an AI co-scientist.
@@ 2 Related Work
Agents for scientific discovery. The increasing accessibility of LLMs and open-source frameworks to integrate them with external environments and custom tools (chase2022langchain; narayanan2025aviary) has spawned multiple agents to either assist in or automate scientific discovery. Some examples include Robin (ghareeb2025robin), which can repurpose existing drugs as potential candidate therapeutics for other diseases, and SciAgents (ghafarollahi2025sciagents), which can traverse knowledge graphs to propose hypotheses about previously unrecognized relationships between scientific concepts. In computational chemistry, ChemCrow (m2024augmenting) and LLM-RDF (ruan2024automatic) can synthesize new compounds and guide end-to-end chemical synthesis. Cactus (mcnaughton2024cactus) can tackle drug and molecular property prediction tasks with domain-specific tools. Further, AstroAgents (saeedi2025astroagents) can process mass spectrometry data to generate plausible hypotheses in the context of existing astrobiology literature.
@@
Besides these domains, a number of agentic workflows have been recently proposed for single-cell biology at different levels of specificity. BioDiscoveryAgent (roohani2025biodiscoveryagent) is designed to propose gene perturbation panels to achieve a target cell phenotype. CellAgent (xiao2024cellagent) involves a planner and executor to generate code blocks, and an evaluator to evaluate the output for three well-defined tasks. However, an agent as a co-scientist should build end-to-end workflows supporting the generation of novel open-ended hypotheses. To this end, CellVoyager (alber2025cellvoyager), BixBench-Agent (mitchener2025bixbench), and Biomni-A1 (huang2025biomni) plan autonomous workflows to derive data-driven insights. The key differences include the environment construction and agentic architecture. However, such agents are currently evaluated on more straightforward scientific reasoning and computational questions, rather than their intended use case.
@@
Benchmarking agents for data-driven scientific discovery. An increase in agents proposed across domains necessitates suitable benchmarks to evaluate the progress. Existing benchmarks for scientific agents share some common characteristics. First, several benchmarks assess scientific thinking in a standalone manner. CORE-Bench (siegel2024corebench) creates tasks to test reproducibility in code repositories. LAB-Bench (laurent2024lab) and HLE (Biomedicine) (phan2025humanity) tests general biological knowledge and reasoning. SciCode (tian2024scicode) evaluates code generation from scientific concepts. Yet none of them pose tasks suitable for evaluating co-scientists. Second, data-driven single-cell benchmarks (majumder2025discoverybench; mitchener2025bixbench; chen2025scienceagentbench) primarily involve statistical or computational problems, and others, such as BaisBench (luo2025benchmarking), are limited to multiple-choice questions alone. As a result, evaluation is constrained to matching against fixed answer sets, which does not reflect the open-ended, exploratory nature of scientific discovery. In addition, BaisBench generates research questions automatically using a single LLM without human oversight, reducing its reliability. gu2024blade also creates multiple-choice questions only, but improves reliability by crowdsourcing annotations of intermediate analysis decisions, and evaluates the open-endedness of scientific discovery by matching these decisions. As a result, the evaluation is constrained to intermediate human preferences and ignores the final agent interpretation. In contrast, our HeurekaBench framework utilizes a multi-LLM pipeline where question generation is divided into suitable subtasks with human supervision. Furthermore, we create both open-ended and multiple-choice research questions which require agents to plan workflows that involve multiple computational steps and conclude with an open-ended, data-driven interpretation, which is evaluated against ground-truth answers from scientific findings.
@@ 3.2 HeurekaBench Creation
The output of this stage is a pool of validated insights. Thus, our framework filters unverifiable insights, directly addressing the reliability gap of co-scientist benchmarks that rely solely on single LLM capabilities (luo2025benchmarking).
@@ 3.3 Evaluation
Evaluating agents on our benchmark poses distinct challenges for both question types. For open-ended research questions, agent analyses may uncover additional conclusions beyond the annotated ground truth. Moreover, in natural-language responses, an agent could potentially rely on prior knowledge from the literature rather than exploring the dataset. Therefore, evaluation of OEQs must go beyond surface-level matching. We adopt G-Eval (liu2023geval) with GPT-4o as the LLM-Judge, assigning ratings between 1 and 5. To ensure scientific rigour, we instruct the judge to first decompose both the response and the ground truth into atomic facts (e.g., conditions, trends, conclusions) and then assess overlap across complete, partial, and missing facts (the entire rubric is provided in Appendix B). An agent receives the highest score only if all ground-truth facts are present and no contradictions occur, while additional non-conflicting findings are not penalized. An illustration of agent evaluation on open-ended questions is presented in Figure 1(c).
@@ 3.4 Instantiating HeurekaBench in single-cell biology
We instantiate the HeurekaBench framework in single-cell biology, particularly for studies on the analysis of scRNA-sequencing datasets (tang2009mrna). First, we curate a pool of $22$ papers published in Nature and Cell journals in 2024 and 2025, with corresponding open-source code repositories and open-access datasets from the CellxGene (czi2025cz) or publication resources. Our choice to restrict to recent publications partially mitigates the risk that agents can rely solely on memorized knowledge. We then applied our insight generation pipeline, utilizing GPT-4o (achiam2023gpt) in InsightExtractor and Claude-4-Sonnet (anthropic2025claude4) in the remaining code modules, to produce $10$ candidate insights per paper and retained only those that could be validated. This process yielded a final pool of $41$ validated insights across $13$ papers (nine from Nature and four from Cell; listed in Appendix C.1). We treat an insight as validated only if the workflow output reproduces the results reported in the paper. While other generated insights may be validated with additional information, we treat them as invalidated within our framework. We provide additional discussion about tasks, publications, and invalidated insights in Appendix C.2, C.3, and C.4 respectively. Finally, using our question generation pipeline, we derived $50$ OEQs and $50$ MCQs from the validated insights, constructing the sc-HeurekaBench, the single-cell biology instantiation of our framework.
@@
During manual workflow review, we observed multiple insights that relied on domain-specific tools and databases (e.g., SCENIC (aibar2017scenic), CellPhoneDB (efremova2020cellphonedb), CellChat (jin2021inference)) as well as machine learning methods (e.g., Non-negative Matrix Factorization). These insights could not be validated because the CodeGenerator hallucinated the usage of these tools or their outputs. Nevertheless, they can serve as a specialized benchmark for evaluating agents’ ability to leverage domain-specific tools, particularly when selected from papers containing other validated insights. We created $12$ OEQs from our collection of $13$ papers that included at least one validated insight, excluding questions that relied on imaging or spatial transcriptomics tools. These questions constitute the sc-HeurekaBench-ToolUsage (sc-HeurekaBench-TU) benchmark, suitable for evaluating agents with access to domain-specific tools (huang2025biomni).
@@ 4.1 Evaluation of Insight Construction
The insight construction stage within the HeurekaBench framework comprises four modules, as described in Section 3.2. To evaluate the InsightExtractor module, we assemble pairs of open-access publications linked with expert findings from two resources. Specifically, we leverage FlyBase (ozturk2024flybase), an openly accessible genome database of Drosophila. We focus on $10$ random gene identifiers to scrape $50$ pairs of publications and corresponding expert findings. The list of genes and other collection details is provided in Appendix D.1. In addition to FlyBase, we also repurpose BixBench (mitchener2025bixbench) to obtain a list of $21$ publications and expert hypotheses pairs. We run our InsightExtractor module on each paper and use GPT-4o as a judge to label if our generated insights are strongly related, weakly related, or unrelated to the expert insight. The instructions for the judge are available in Appendix D.2. We show our results in Figure 2(a). For FlyBase, we obtain $44$ strongly, two weakly, and four unrelated findings. At the same time, for BixBench, we have $14$ strongly, four weakly, and three unrelated. Hence, InsightExtractor module is capable of determining expert-level insights from scientific publications.
Furthermore, we collect pairs of expert-created insights and corresponding multi-step workflows from InsightBench (sahu2025insightbench) to assess our CodeDescriber and CodeMatcher modules. We convert each step within a workflow into a code file and create a code repository with these files. We restrict the repository to $50$ insights (and $215$ code files) to match the maximum size of a real-world repository from our set of validated publications in the sc-HeurekaBench  [^1] Out of the initial $22$ papers, $18$ have less than $100$ code files. The CodeDescriber annotates the code scripts, and given the insight description, CodeMatcher retrieves the relevant files. Although this is a challenging setting, out of $215$ scripts, $158$ scripts were correctly matched to the respective insights, and an average of $\mathbf{74.6\%}$ of files are retrieved correctly across all insights. Additionally, we also plot the number of insights per incorrectly retrieved files in Figure 2(b), which shows that our modules can select the relevant files corresponding to insights for code generation.
@@ 4.2 Evaluation of Single-Cell Biology agents
We benchmark three state-of-the-art agents for single-cell biology on the sc-HeurekaBench: Biomni (huang2025biomni), CellVoyager (alber2025cellvoyager), and BixBench-Agent (mitchener2025bixbench). Biomni (huang2025biomni) comprises Biomni-E1, an extensive biomedical environment that contains $105$ software packages, $59$ databases, and $150$ specialized biomedical tools, including cell biology and genomics domains, and Biomni-A1, an agent that navigates Biomni-E1 to build autonomous multi-step workflows for data analysis on experimental data. CellVoyager (alber2025cellvoyager) is designed to propose novel hypotheses from existing studies and validate them. However, unlike Biomni, it follows a rigid architecture, where each step contains specific LLMs for sequentially planning, coding, handling feedback, interpreting outputs, and summarizing. Consequently, we remove the hypothesis proposal step and instruct to answer research questions. Finally, BixBench-Agent (mitchener2025bixbench) is a more black-box agent that relies on the Aviary custom environment (narayanan2025aviary), with predefined bioinformatics packages.
@@ 4.3.1 Planner ablations
Amongst all LLMs, Claude-4-Sonnet (anthropic2025claude4) achieves the highest overall performance across OEQs and MCQs. Particularly in OEQs, it outperforms the second-best model with a significant margin ($2.58$ against $2.08$), highlighting the benefits of frontier closed-source models. Within the Qwen (yang2025qwen3) model family, performance consistently improves with increasing model size, which suggests that model parameter scale is crucial, and the thinking variant provides additional gains in correctness on OEQs due to its better reasoning abilities (+0.28+0.28 from non-thinking and +0.38+0.38 from smaller model). Among open-source models, the best-performing model was GPT-OSS-120B (agarwal2025gpt), a recent OpenAI model designed for use within agentic workflows. Overall, these ablations confirm that agent performance as a co-scientist is highly dependent on model family, scale, and reasoning style.
@@
We perform an alignment study between the correctness scores assigned by three different closed-source LLMs (including the proposed GPT-4o) as judges. In particular, we select Claude-4.5-Sonnet (anthropic2025claude45) and Gemini-2.5-Pro (comanici2025gemini), which are not used as planner LLMs within the Biomni agent and belong to a different model series than GPT. We consider the three best-performing planner models, i.e., Claude-4-Sonnet, GPT-OSS-120B, and Qwen3-235B-thinking and show the ranks assigned by the three judges in Figure 3. We observe that all judges agree on the ranking of the planner models. Next, for each model, we compute more fine-grained inter-rater agreement metrics comparing the correctness scores assigned by GPT-4o and other judges. In particular, we report Spearman’s rank correlation and unweighted Cohen’s kappa ($\kappa$) with quadratic penalty, which are suitable for ordinal discrete ratings. We find that the mean correlation across three planner models between GPT-4o and Claude-4.5-Sonnet judges is $0.84_{\pm 0.03}$, whereas it is $0.79_{\pm 0.01}$ with Gemini-2.5-Pro. Similarly, the $\kappa$ is $0.81_{\pm 0.03}$ and $0.71_{\pm 0.04}$ respectively. This study indicates high agreement between different LLM judges. We provide detailed values of these metrics in Appendix D.4 and use GPT-4o for all subsequent evaluations.

### v2

@@ 1 Introduction
Post-training algorithms for large language models (LLMs) with reinforcement learning (RL) have led to rapid progress in improving their reasoning capabilities (Rafailov et al., 2023; Shao et al., 2024; Chai et al., 2025). As a result, LLM-based agents have been developed to solve complex tasks through multi-turn interactions with an external environment (Huang et al., 2022; Yao et al., 2023; Yang et al., 2024). These systems can be broadly characterized by a loop consisting of three stages: in the think stage, the agent needs to understand the user request and create a plan; in the act stage it generates an action depending on the plan; and in the observe stage it updates the context with the feedback of action execution for the next cycle.
Agentic systems offer promise for advancing scientific discovery by actively generating new hypotheses, designing and evaluating experiments, and iteratively improving through feedback (M. Bran et al., 2024; Ghareeb et al., 2025; Saeedi et al., 2025; Huang et al., 2025a). In this capacity, they could act as AI co-scientists, complementing human expertise and accelerating the scientific discovery process (Yamada et al., 2025; Gottweis et al., 2025). One scientific domain that has recently seen an influx of agents is single-cell biology (Xiao et al., 2024; Roohani et al., 2025; Alber et al., 2025; Mitchener et al., 2025; Huang et al., 2025b). A typical real-world use case in this domain involves autonomously proposing novel hypotheses, performing complex analyses of single-cell RNA-seq experimental datasets and delivering scientifically meaningful findings.
@@
Most existing data-driven scientific discovery benchmarks test task instruction following and answering a single computational question from experimental data (e.g., How many miRNAs remain significant at $p\leq 0.05$ after Benjamini-Hochberg correction? from Mitchener et al. (2025) and Train a VAE model and perform a 1-vs-all differential expression test for each cell type from Chen et al. (2025)). However, a co-scientist should autonomously plan these questions as sub-steps within the workflow, rather than explicit user instructions. The closest to our work is BaisBench (Luo et al., 2025); however, it relies solely on the capabilities of LLMs to formulate questions from scientific studies, leading to invalidated questions that may not be possible to answer accurately. Thus, existing benchmarks lack an appropriate framework for evaluating open-ended, free-form, data-driven scientific responses that are required for an AI co-scientist.
@@ 2 Related Work
Agents for scientific discovery. The increasing accessibility of LLMs and open-source frameworks to integrate them with external environments and custom tools (Chase, 2022; Narayanan et al., 2025) has spawned multiple agents to either assist in or automate scientific discovery. Some examples include Robin (Ghareeb et al., 2025), which can repurpose existing drugs as potential candidate therapeutics for other diseases, and SciAgents (Ghafarollahi and Buehler, 2025), which can traverse knowledge graphs to propose hypotheses about previously unrecognized relationships between scientific concepts. In computational chemistry, ChemCrow (M. Bran et al., 2024) and LLM-RDF (Ruan et al., 2024) can synthesize new compounds and guide end-to-end chemical synthesis. Cactus (McNaughton et al., 2024) can tackle drug and molecular property prediction tasks with domain-specific tools. Further, AstroAgents (Saeedi et al., 2025) can process mass spectrometry data to generate plausible hypotheses in the context of existing astrobiology literature.
@@
Besides these domains, a number of agentic workflows have been recently proposed for single-cell biology at different levels of specificity. BioDiscoveryAgent (Roohani et al., 2025) is designed to propose gene perturbation panels to achieve a target cell phenotype. CellAgent (Xiao et al., 2024) involves a planner and executor to generate code blocks, and an evaluator to evaluate the output for three well-defined tasks. However, an agent as a co-scientist should build end-to-end workflows supporting the generation of novel open-ended hypotheses. To this end, CellVoyager (Alber et al., 2025), BixBench-Agent (Mitchener et al., 2025), and Biomni-A1 (Huang et al., 2025b) plan autonomous workflows to derive data-driven insights. The key differences include the environment construction and agentic architecture. However, such agents are currently evaluated on more straightforward scientific reasoning and computational questions, rather than their intended use case.
@@
Benchmarking agents for data-driven scientific discovery. An increase in agents proposed across domains necessitates suitable benchmarks to evaluate the progress. Existing benchmarks for scientific agents share some common characteristics. First, several benchmarks assess scientific thinking in a standalone manner. CORE-Bench (Siegel et al., 2024) creates tasks to test reproducibility in code repositories. LAB-Bench (Laurent et al., 2024) and HLE (Biomedicine) (Phan et al., 2025) tests general biological knowledge and reasoning. SciCode (Tian et al., 2024) evaluates code generation from scientific concepts. Yet none of them pose tasks suitable for evaluating co-scientists. Second, data-driven single-cell benchmarks (Majumder et al., 2025; Mitchener et al., 2025; Chen et al., 2025) primarily involve statistical or computational problems, and others, such as BaisBench (Luo et al., 2025), are limited to multiple-choice questions alone. As a result, evaluation is constrained to matching against fixed answer sets, which does not reflect the open-ended, exploratory nature of scientific discovery. In addition, BaisBench generates research questions automatically using a single LLM without human oversight, reducing its reliability. Gu et al. (2024a) also creates multiple-choice questions only, but improves reliability by crowdsourcing annotations of intermediate analysis decisions, and evaluates the open-endedness of scientific discovery by matching these decisions. As a result, the evaluation is constrained to intermediate human preferences and ignores the final agent interpretation. In contrast, our HeurekaBench framework utilizes a multi-LLM pipeline where question generation is divided into suitable subtasks with human supervision. Furthermore, we create both open-ended and multiple-choice research questions which require agents to plan workflows that involve multiple computational steps and conclude with an open-ended, data-driven interpretation, which is evaluated against ground-truth answers from scientific findings.
@@ 3.2 HeurekaBench Creation
The output of this stage is a pool of validated insights. Thus, our framework filters unverifiable insights, directly addressing the reliability gap of co-scientist benchmarks that rely solely on single LLM capabilities (Luo et al., 2025).
@@ 3.3 Evaluation
Evaluating agents on our benchmark poses distinct challenges for both question types. For open-ended research questions, agent analyses may uncover additional conclusions beyond the annotated ground truth. Moreover, in natural-language responses, an agent could potentially rely on prior knowledge from the literature rather than exploring the dataset. Therefore, evaluation of OEQs must go beyond surface-level matching. We adopt G-Eval (Liu et al., 2023) with GPT-4o as the LLM-Judge, assigning ratings between 1 and 5. To ensure scientific rigour, we instruct the judge to first decompose both the response and the ground truth into atomic facts (e.g., conditions, trends, conclusions) and then assess overlap across complete, partial, and missing facts (the entire rubric is provided in Appendix B). An agent receives the highest score only if all ground-truth facts are present and no contradictions occur, while additional non-conflicting findings are not penalized. An illustration of agent evaluation on open-ended questions is presented in Figure 1(c).
@@ 3.4 Instantiating HeurekaBench in single-cell biology
We instantiate the HeurekaBench framework in single-cell biology, particularly for studies on the analysis of scRNA-sequencing datasets (Tang et al., 2009). First, we curate a pool of $22$ papers published in Nature and Cell journals in 2024 and 2025, with corresponding open-source code repositories and open-access datasets from the CellxGene (CZI Cell Science Program et al., 2025) or publication resources. Our choice to restrict to recent publications partially mitigates the risk that agents can rely solely on memorized knowledge. We then applied our insight generation pipeline, utilizing GPT-4o (Achiam et al., 2023) in InsightExtractor and Claude-4-Sonnet (Anthropic, 2025b) in the remaining code modules, to produce $10$ candidate insights per paper and retained only those that could be validated. This process yielded a final pool of $41$ validated insights across $13$ papers (nine from Nature and four from Cell; listed in Appendix C.1). We treat an insight as validated only if the workflow output reproduces the results reported in the paper. While other generated insights may be validated with additional information, we treat them as invalidated within our framework. We provide additional discussion about tasks, publications, and invalidated insights in Appendix C.2, C.3, and C.4 respectively. Finally, using our question generation pipeline, we derived $50$ OEQs and $50$ MCQs from the validated insights, constructing the sc-HeurekaBench, the single-cell biology instantiation of our framework.
@@
During manual workflow review, we observed multiple insights that relied on domain-specific tools and databases (e.g., SCENIC (Aibar et al., 2017), CellPhoneDB (Efremova et al., 2020), CellChat (Jin et al., 2021)) as well as machine learning methods (e.g., Non-negative Matrix Factorization). These insights could not be validated because the CodeGenerator hallucinated the usage of these tools or their outputs. Nevertheless, they can serve as a specialized benchmark for evaluating agents’ ability to leverage domain-specific tools, particularly when selected from papers containing other validated insights. We created $12$ OEQs from our collection of $13$ papers that included at least one validated insight, excluding questions that relied on imaging or spatial transcriptomics tools. These questions constitute the sc-HeurekaBench-ToolUsage (sc-HeurekaBench-TU) benchmark, suitable for evaluating agents with access to domain-specific tools (Huang et al., 2025b).
@@ 4.1 Evaluation of Insight Construction
The insight construction stage within the HeurekaBench framework comprises four modules, as described in Section 3.2. To evaluate the InsightExtractor module, we assemble pairs of open-access publications linked with expert findings from two resources. Specifically, we leverage FlyBase (Öztürk-Çolak et al., 2024), an openly accessible genome database of Drosophila. We focus on $10$ random gene identifiers to scrape $50$ pairs of publications and corresponding expert findings. The list of genes and other collection details is provided in Appendix D.1. In addition to FlyBase, we also repurpose BixBench (Mitchener et al., 2025) to obtain a list of $21$ publications and expert hypotheses pairs. We run our InsightExtractor module on each paper and use GPT-4o as a judge to label if our generated insights are strongly related, weakly related, or unrelated to the expert insight. The instructions for the judge are available in Appendix D.2. We show our results in Figure 2(a). For FlyBase, we obtain $44$ strongly, two weakly, and four unrelated findings. At the same time, for BixBench, we have $14$ strongly, four weakly, and three unrelated. Hence, InsightExtractor module is capable of determining expert-level insights from scientific publications.
Furthermore, we collect pairs of expert-created insights and corresponding multi-step workflows from InsightBench (Sahu et al., 2025) to assess our CodeDescriber and CodeMatcher modules. We convert each step within a workflow into a code file and create a code repository with these files. We restrict the repository to $50$ insights (and $215$ code files) to match the maximum size of a real-world repository from our set of validated publications in the sc-HeurekaBench  [^1] Out of the initial $22$ papers, $18$ have less than $100$ code files. The CodeDescriber annotates the code scripts, and given the insight description, CodeMatcher retrieves the relevant files. Although this is a challenging setting, out of $215$ scripts, $158$ scripts were correctly matched to the respective insights, and an average of $\mathbf{74.6\%}$ of files are retrieved correctly across all insights. Additionally, we also plot the number of insights per incorrectly retrieved files in Figure 2(b), which shows that our modules can select the relevant files corresponding to insights for code generation.
@@ 4.2 Evaluation of Single-Cell Biology agents
We benchmark three state-of-the-art agents for single-cell biology on the sc-HeurekaBench: Biomni (Huang et al., 2025b), CellVoyager (Alber et al., 2025), and BixBench-Agent (Mitchener et al., 2025). Biomni (Huang et al., 2025b) comprises Biomni-E1, an extensive biomedical environment that contains $105$ software packages, $59$ databases, and $150$ specialized biomedical tools, including cell biology and genomics domains, and Biomni-A1, an agent that navigates Biomni-E1 to build autonomous multi-step workflows for data analysis on experimental data. CellVoyager (Alber et al., 2025) is designed to propose novel hypotheses from existing studies and validate them. However, unlike Biomni, it follows a rigid architecture, where each step contains specific LLMs for sequentially planning, coding, handling feedback, interpreting outputs, and summarizing. Consequently, we remove the hypothesis proposal step and instruct to answer research questions. Finally, BixBench-Agent (Mitchener et al., 2025) is a more black-box agent that relies on the Aviary custom environment (Narayanan et al., 2025), with predefined bioinformatics packages.
@@ 4.3.1 Planner ablations
Amongst all LLMs, Claude-4-Sonnet (Anthropic, 2025b) achieves the highest overall performance across OEQs and MCQs. Particularly in OEQs, it outperforms the second-best model with a significant margin ($2.58$ against $2.08$), highlighting the benefits of frontier closed-source models. Within the Qwen (Yang et al., 2025) model family, performance consistently improves with increasing model size, which suggests that model parameter scale is crucial, and the thinking variant provides additional gains in correctness on OEQs due to its better reasoning abilities (+0.28+0.28 from non-thinking and +0.38+0.38 from smaller model). Among open-source models, the best-performing model was GPT-OSS-120B (Agarwal et al., 2025), a recent OpenAI model designed for use within agentic workflows. Overall, these ablations confirm that agent performance as a co-scientist is highly dependent on model family, scale, and reasoning style.
@@
We perform an alignment study between the correctness scores assigned by three different closed-source LLMs (including the proposed GPT-4o) as judges. In particular, we select Claude-4.5-Sonnet (Anthropic, 2025a) and Gemini-2.5-Pro (Comanici et al., 2025), which are not used as planner LLMs within the Biomni agent and belong to a different model series than GPT. We consider the three best-performing planner models, i.e., Claude-4-Sonnet, GPT-OSS-120B, and Qwen3-235B-thinking and show the ranks assigned by the three judges in Figure 3. We observe that all judges agree on the ranking of the planner models. Next, for each model, we compute more fine-grained inter-rater agreement metrics comparing the correctness scores assigned by GPT-4o and other judges. In particular, we report Spearman’s rank correlation and unweighted Cohen’s kappa ($\kappa$) with quadratic penalty, which are suitable for ordinal discrete ratings. We find that the mean correlation across three planner models between GPT-4o and Claude-4.5-Sonnet judges is $0.84_{\pm 0.03}$, whereas it is $0.79_{\pm 0.01}$ with Gemini-2.5-Pro. Similarly, the $\kappa$ is $0.81_{\pm 0.03}$ and $0.71_{\pm 0.04}$ respectively. This study indicates high agreement between different LLM judges. We provide detailed values of these metrics in Appendix D.4 and use GPT-4o for all subsequent evaluations.